# Competition of Mechanisms: Tracing How Language Models Handle Facts and Counterfactuals

### v1

@@ Methods.
To understand the main source of factual and counterfactual information introduced in the last position, we apply the logit inspection method to the output of the attention and MLP blocks before they are added to the residual stream. We analyze the values of $t_{\mathrm{fact}}$ and $t_{\mathrm{cofa}}$ in the the logit distribution at $W_{U}\mathbf{a_{N}}^{l}$ and $W_{U}\mathbf{m_{N}}^{l}$, where $N$ denotes the last token in the sequence. A positive value of $\Delta_{\mathrm{cofa}}=\text{Logit}(t_{\mathrm{cofa}})-\text{Logit}(t_{\mathrm{fact}})$ at $W_{U}\mathbf{a_{N}}^{l}$ (resp. $W_{U}\mathbf{m_{N}}^{l}$) means that the attention map (resp. MLP) is increasing the likelihood of the counterfactual attribute in the last position at layer $l$ relative to the factual one.
@@ Results.
Figure 3 compares the $\Delta_{\mathrm{cofa}}$ values added to the residual stream by the attention and MLP blocks at all the GPT-2-small layers averaged on the whole dataset. The left panel shows the contribution to $\Delta_{\mathrm{cofa}}$ of the attention blocks summing over all the attention heads. Until layer 4, $\Delta_{\mathrm{cofa}}$ is approximately zero, meaning that the attention maps do not add information about either $t_{\mathrm{fact}}$ or $t_{\mathrm{cofa}}$. From layers 5 to 10, the value of $\Delta_{\mathrm{cofa}}$ increases, reaching a peak of about 1.4 at layer 9. The last layer is the only one that favors $t_{\mathrm{fact}}$ with a sizeable negative value of $\Delta_{\mathrm{cofa}}$ around -0.8. In the next section, we will see that the decrease of $\Delta_{\mathrm{cofa}}$ in the last two layers is mainly due to the contribution of two attention heads, L10H7 and L11H10, which explain most of the total negative contribution to $\Delta_{\mathrm{cofa}}$.
The right panel shows the evolution of $\Delta_{\mathrm{cofa}}$ for the MLP blocks. Similar to the trend observed for the attention blocks, this contribution remains close to zero in the early layers of the networks and grows only after layer 8. The total contribution of the MLPs to $\Delta_{\mathrm{cofa}}$ is positive but much smaller than the one given by the attention blocks and significantly different from zero only at layers 9 and 10, where it reaches its maximum value around 0.75. Attention blocks and MLPs behave consistently across the network, favoring the counterfactual attribute in all the layers except for the attention contribution in the last layer. The attention blocks are in charge of adding most of the information about $t_{\mathrm{cofa}}$ to the residual stream of the last position, consistently with the findings of Geva et al. (2023). The next section delves into the analysis of the attention heads to better understand their behavior. In particular, we will investigate if some heads have a special role in favoring factual or counterfactual tokens and identify the position from which they primarily extract information.
6.3 Effects of Attention Heads
In this section, we study the role played by each attention head in promoting the factual vs counterfactual token. We will find that the behavior of the attention heads of the final layers is highly specialized, and the inspection of the heads that contribute most to $\Delta_{\mathrm{cofa}}$ allows finding from which position most of the information is brought to the final one. We finally show that we can significantly increase the factual predictions by simply up-weighting the few specialized attention weights where most factual information flows.
Methods.
We analyze the attention heads separately with the logit inspection method by projecting the outputs of each attention head to the last sequence position $N$ in the vocabulary space. More precisely, we consider $\Delta_{\mathrm{cofa}}=\text{Logit}(t_{\mathrm{cofa}})-\text{Logit}(t_{\mathrm{fact}})$ where the logits are taken from the projection $W_{U}\mathbf{a_{N}}^{h,l}$ for each head $h$.
@@
Figure 4 (left) shows on the $y$-axis the individual contribution of each head to $\Delta_{\mathrm{cofa}}$ averaged over the dataset. On the $x$-axis, we plot the layers of the network. A blue cell indicates a positive $\Delta_{\mathrm{cofa}}$ or that a given head is favoring $t_{\mathrm{cofa}}$. Conversely, a red cell indicates a negative $\Delta_{\mathrm{cofa}}$ or that a given head is favoring $t_{\mathrm{fact}}$. Figure 4 (left) shows that few heads are responsible for promoting or suppressing the copy of $t_{\mathrm{cofa}}$ to the last position. For example, the sum of L7H2 and L7H10 equals 75% of the large positive $\Delta_{\mathrm{cofa}}$ contribution of layer 7. The sum of L9H6 and L9H9 explains 65% of the $\Delta_{\mathrm{cofa}}$ at layer 9. In layers 10 and 11, the decrease of $\Delta_{\mathrm{cofa}}$ shown in Fig. 3 (left) is almost entirely due to the action of two heads: L10H7 and L11H10, which alone explain almost the 70% of the total negative contribution to $\Delta_{\mathrm{cofa}}$ in the entire network (33% and 37% respectively). McDougall et al. (2023) showed that these two heads are responsible for suppressing the copy mechanisms in GPT-2-small. In our setting, the joint ablation of these two heads decreases the factual recall of GPT-2-small from 4.13% to 0.65%.
Most interestingly, all highly activated heads attend to the same position. Figure 4 (right) shows the last rows of the most activated attention maps (we report the full maps in Figure S6 of the Appendix). Even though each attention head may have a different role in supporting either $t_{\mathrm{fact}}$ or $t_{\mathrm{cofa}}$, they all strongly attend to the attribute position. This is expected for the heads that support the copy mechanism since they read from the sequence position where $t_{\mathrm{cofa}}$ is better represented (see Sec. 6.1), while it is more surprising for the heads that favor the factual token. Indeed, an attention head can favor a factual response by increasing the logit of $t_{\mathrm{fact}}$ or decreasing the logit of $t_{\mathrm{cofa}}$ or both. In this case, L10H7 and L11H10 read from the attribute position because it is easier to give a lower value to the logit of $t_{\mathrm{cofa}}$ without at the same time penalizing too much $t_{\mathrm{fact}}$. In these two heads, the logit of $t_{\mathrm{fact}}$ written to the last sequence position is smaller than the mean of the two layers, but the logit $t_{\mathrm{cofa}}$, -1.13 for L10H7 and -1.05 for L11H10, are by far the lowest of all the heads in the network. This picture also holds true for larger GPT-2 transformers and Pythia: in all cases, the most activated heads have a strongly activated attention edge between the attribute and the last position, including those favoring $t_{\mathrm{fact}}$ (see Fig. S3 of the Appendix).
6.4 Improving the Factual Recall
@@
Because the factual mechanism is promoted by only a few strongly activated heads (L10H7 and L11H10 in GPT-2, and mostly L17H28, L20H18, and L21H8 in Pythia, see Fig. S4) and because most of the information flows from the attribute position (see Fig. 4-right and Sec. 6.2), enhancing the weight of this path will likely improve the factual recall of the model. In this section, we show that enhancing the value of a few well-localized attention edges can largely improve the factual recall of the model.
Methods.
To this end, we multiply by a common factor $\alpha$ the attribute position in the last row of L10H7 and L11H10 for GPT-2, and L17H28, L20H18, and L21H8 in Pythia (see Appendix, Fig. S4). We search the optimal value of $\alpha$ on the grid $[2,5,10,100]$.
Figure 5: Boosting factual recall with attention score modification. The figure shows the increase in factual recall for GPT-2 and Pythia-6.9b, achieved by increasing the attention score of the attribute position in the last row of the attention map by a factor of 5 for the L10H7 and L11H10 in GPT-2 and L17H28, 498 L20H18, and L21H8 in Pythia.
@@ Results.
Figure 5 shows the gain of factual recall obtained using $\alpha=5$ for both GPT-2 for Pythia. The factual error rate decreases from 96% to 50% in GPT-2 and from 70% to 50% in Pythia. This result is remarkable since we modify only two entries in the attention map out of the 117M parameters of GPT-2 and three entries out of the 6.9B parameters of Pythia. This highlights the importance of the analysis of Sec. 6.3 and Sec. 6.2 in finding the detailed role played by the individual units of each transformer layer.
@@ 6.5 What Word Choices Intensify the Competition?
We next focus on understanding how the similarity between $t_{\mathrm{fact}}$ and $t_{\mathrm{cofa}}$ in our dataset affects the mechanism described in the previous sections.
Methods.
We measure the similarity of the attributes using a pre-trained Word2Vec model (Mikolov et al., 2013) from Gensim Library (Řehůřek and Sojka, 2010), which generates 300-dimensional word vectors.
Figure 6: Prediction Frequency of Factual Token by Similarity Level. We show the percentage of $t_{\mathrm{fact}}$ predictions within each bin compared to the entire dataset (represented by a dotted line) across various model sizes. Each bin has a different range of cosine similarity between $t_{\mathrm{fact}}$ and $t_{\mathrm{cofa}}$. We can notice that more similar $t_{\mathrm{fact}}$ and $t_{\mathrm{cofa}}$ are, and the factual mechanism is stronger.
@@ Results.
We divide the dataset into ten equally sized subsets according to the cosine similarity between the vectors representing $t_{\mathrm{fact}}$ and $t_{\mathrm{cofa}}$. We then measure the frequency of predicting $t_{\mathrm{fact}}$ within each subset.
Figure 6 shows the outcomes across various model sizes. We can see that the factual recall is higher when $t_{\mathrm{fact}}$ and $t_{\mathrm{cofa}}$ are more similar. This trend is observed across all model sizes, with larger models generally exhibiting a stronger correlation and a higher overall propensity for predicting the factual token. This suggests that larger models are more sensitive to the similarity between $t_{\mathrm{cofa}}$ and $t_{\mathrm{fact}}$ and are more likely to rely on factual recall, even when the similarity is not exceptionally high. This finding aligns with the notion that larger models have a greater capacity to store and retrieve factual information.
@@ 7 Discussion and Future Work
To this point, we have analyzed in detail the mechanism by which LLMs handle the competition between facts learned during training and counterfactual information found in-context. Our findings are consistent with Geva et al. (2023), showing that late attention blocks write most of the information to the last layer when adding a counterfactual premise. Surprisingly, however, we found that the largest contribution to the factual prediction of the network mostly comes from the suppression of the counterfactual token read from the attribute position rather than the promotion of the factual token from the subject position. Consistently with McDougall et al. (2023), we found that few highly specialized heads suppress the counterfactual information. Moreover, we also found that up-weighting only two or three attention entries of these heads is enough to increase substantially the number of factual responses of the network. With an approach similar to ours, Yu et al. (2023) found that more heads can promote the factual mechanism, also in early layers, but found it challenging to improve the factual responses by scaling up the weights of the attention maps. This discrepancy can be due to the broader set of topics we include in our prompts, which allowed us to select fewer, more specialized heads, to the different ways the prompts are framed, or also to our more focused modification of the attention maps. As future research directions, we aim to analyze more in-depth how our findings depend on the prompt structure and whether the promotion of factual responses by suppressing the counterfactuals generalizes to larger models and a more comprehensive variety of datasets.
@@ 8 Conclusion
In this work, we have proposed the formulation of the competition of algorithms as a powerful interpretation when LLMs need to handle multiple mechanisms, only one of which leads to the correct answer. We deployed two mechanistic interpretability tools, logit inspection and attention modification, and identified critical positions and model components involved in competing for the mechanisms. Finally, we discovered a few localized positions in the attention map, which largely control the strength of the factual mechanism. Our study sheds light on future work on interpretability research for LLMs.
@@
Our focus was mainly on GPT-2 models, particularly GPT-2-small, which possess considerably fewer parameters than current state-of-the-art models. Furthermore, our experiments and insights are heavily grounded in the interpretability within the embedding space of the model’s inner components. This approach is reliable and extensively employed in mechanistic interpretability research (Dar et al., 2023; Geva et al., 2022; Halawi et al., 2023). The logit inspection method, although commonly employed in previous work, can occasionally fail to reflect the actual importance of some vocabulary items, especially in the early layers of the network (Belrose et al., 2023). Finally, our research notes that the competitive dynamics intensify with increasing model size. However, as Wei et al. (2023) points out, this trend may exhibit a U-shaped curve, decreasing beyond a specific model size threshold.
@@ Acknowledgment
This material is based in part upon works supported by the German Federal Ministry of Education and Research (BMBF): Tübingen AI Center, FKZ: 01IS18039B; by the Machine Learning Cluster of Excellence, EXC number 2064/1 – Project number 390727645; by the John Templeton Foundation (grant #61156); by a Responsible AI grant by the Haslerstiftung; and an ETH Grant (ETH-19 21-1). Alberto Cazzaniga and Diego Doimo are supported by the project “Supporto alla diagnosi di malattie rare tramite l’intelligenza artificiale" CUP: F53C22001770002. Alberto Cazzaniga received funding by the European Union – NextGenerationEU within the project PNRR "PRP@CERIC" IR0000028 - Mission 4 Component 2 Investment 3.1 Action 3.1.1. Zhijing Jin is supported by PhD fellowships from the Future of Life Institute and Open Philanthropy.
@@
Appendix A Data Composition Details
[…]
Appendix B Experimental Details
We deploy the pre-trained GPT-2 and Pythia models from the Huggingface Hub Wolf et al. (2020), and inspect the residual streams by the LogitLens tool in the TransformerLens library (Nanda and Bloom, 2022).
Appendix C Experiments for Pythia-6.9b

### v2

@@ 6.2 Intermediate Inspection of Attention and MLP Blocks
Hence, we inspect the margin of the added logit of $t_{\mathrm{cofa}}$ over that of $t_{\mathrm{fact}}$ in each block, represented by $\Delta_{\mathrm{cofa}}:=\text{BlockLogit}(t_{\mathrm{cofa}})-\text{BlockLogit}(t_{\mathrm{fact}})$. To this end, we apply the logit inspection method to analyze the logit distribution at $W_{U}\mathbf{a_{N}}^{l}$ and $W_{U}\mathbf{m_{N}}^{l}$, where $N$ denotes the last token position in the sequence. The logit contribution of the attention block is the sum over that of all the attention heads. As for the result, a positive value of $\Delta_{\mathrm{cofa}}$ for a block means that it supports the counterfactual mechanism in the competition, and a negative value indicates suppression.
(a) Direct contribution to $\Delta_{\mathrm{cofa}}$ of all heads in GPT-2. Heads favoring $t_{\mathrm{fact}}$ are colored in blue, and those favoring $t_{\mathrm{cofa}}$ in red.
(b) Attention scores of relevant heads for the last position. A large attention score in the attribute position is found in all highly activated heads.
Figure 4: Attention pattern for relevant attention heads.
@@ Results.
We quantify the contribution of each block in each layer by plotting the $\Delta_{\mathrm{cofa}}$ values in Figure 3.
Finding 1: The attention blocks play a larger role in the competition of mechanisms than the MLP blocks. Contrasting the $\Delta_{\mathrm{cofa}}$ margin of the added logits of the attention blocks in Figure 3(a) and MLP blocks in Figure 3(b), we see that the size of $\Delta_{\mathrm{cofa}}$ is almost always larger in the attention blocks than in MLP blocks. This is consistent with the work of Geva et al. (2023) showing that the attention blocks adds most of the information about $t_{\mathrm{cofa}}$ to the residual stream.
Finding 2: Only late but not early layers contribute to the competition of the mechanisms. We find that the early layers have almost no contribution to the competition of the mechanisms, reflected by the close-to-zero margin $\Delta_{\mathrm{cofa}}$ in Layer 0-4 for both types of blocks. However, later layers contribute to substantially to the increase of the margin $\Delta_{\mathrm{cofa}}$, by a relatively smaller rate for the MLP blocks, and a larger overall rate for the attention blocks, together with a large variance.
Note that we observe a negative $\Delta_{\mathrm{cofa}}$ around -0.8 in the last attention block, somewhat favoring $t_{\mathrm{fact}}$, which might be since the factual information is moved to the last position in the last layers, as already noted by Geva et al. 2023.
6.3 Microscopic Inspection of Individual Attention Heads
Beyond the overall contributions of the attention block, we further study the contribution of each individual attention head in this section.
Method.
We analyze the effect of each individual attention head with the logit inspection method by projecting the outputs of each attention head to the last sequence position $N$ in the vocabulary space. Formally, we consider $\Delta_{\mathrm{cofa}}=\text{HeadLogit}(t_{\mathrm{cofa}})-\text{HeadLogit}(t_{\mathrm{fact}})$ with the logits from the projection $W_{U}\mathbf{a_{N}}^{h,l}$ of each head $h$. Here $\mathbf{a_{N}}$ is the output of the attention head $h$ after it has been processed by the output matrix of the Attention Block but before its sum to the residual stream.
@@
We plot the contributions of individual attention heads to $\Delta_{\mathrm{cofa}}$ in Figure 4, and introduce the main findings as follows.
Finding 1: A few specialized attention heads contribute the most to the competition. As we can see from the overall contributions of all attention heads across all the layers in Figure 4(a), several attention heads (e.g., L9H6, L9H9, L10H0, and L10H10) strongly promote the counterfactual mechanism, i.e., with a positive value of $\Delta_{\mathrm{cofa}}$ colored in dark red, and two attention heads (L10H7 and L11H10) strongly support the factual mechanism instead, reflected by the large negative $\Delta_{\mathrm{cofa}}$ in dark blue.
For example, the sum of L7H2 and L7H10 equals 75% of the large positive $\Delta_{\mathrm{cofa}}$ contribution of Layer 7. The sum of L9H6 and L9H9 explains 65% of the $\Delta_{\mathrm{cofa}}$ at Layer 9.
On the other hand, the two attention heads, L10H7 and L11H10, explain almost the 70% of the total negative contribution to $\Delta_{\mathrm{cofa}}$ in the entire network (33% and 37% respectively). This also explains the reason behind the negative $\Delta_{\mathrm{cofa}}$ in Figure 3(a) of the previous section. Our study is consistent with McDougall et al. (2023) showing that these two heads are responsible for suppressing the copy mechanisms in GPT-2 small. In our setting, the joint ablation of these two heads decreases the factual recall of GPT-2 small from 4.13% to 0.65%.
Finding 2: All the highly activated heads attend to the same position – the attribute token. Focusing on the heads with large absolute values of $\Delta_{\mathrm{cofa}}$, we show the attention scores of the last position $N$ to different tokens in Figure 4(b). Expectedly, the major heads supporting the counterfactual mechanism (those in red) attend to the attribute position because they need to copy this token for the prediction, which echoes the findings in Section 6.1.
However, it is surprising to see the other heads supporting the factual mechanism (those in blue) also mainly attend to the counterfactual attribute token. We find that those heads read from the attribute position to give a lower value to the logit of $t_{\mathrm{cofa}}$, which might be an easier operation for it to learn than increasing the logit of the factual token. The evidence is that, in these two heads, the logit of $t_{\mathrm{fact}}$ is smaller than the mean of the two layers, but the logit of $t_{\mathrm{cofa}}$ (which is -1.13 for L10H7 and -1.05 for L11H10) are the lowest of all the heads in the network.
We include supplementary analyses showing the consistency of Finding 2 on Pythia in Appendix A, and provide the full attention maps with attention scores between every pair of tokens for these heads in Appendix Section B.2.
6.4 Intrinsic Intervention by Attention Modification
@@
After understanding where the two mechanisms take place, we use the insights to intervene on the model internal states. Specifically, we perform model editing to alter the factual mechanism, which concentrates on a few strongly activated attention heads (L10H7 and L11H10 in GPT-2, and mostly L17H28, L20H18, and L21H8 in Pythia, see Section A.3), and has most of the information flowing from the attribute position (see Figure 4-right and Section 6.2). In the following, we show that enlarging the value of a few well-localized attention values can largely improve the factual recall of the model.
Method.
We utilize the attention modification method in Eq. 5 to apply a multiplier of $\alpha$ to the attention weights of the last token to the attribute position in L10H7 and L11H10 for GPT-2, and L17H28, L20H18, and L21H8 in Pythia. To choose the $\alpha$ value, we perform a grid search over $[2,5,10,100]$ to maximize the factual recall rate of the model. We find that $\alpha=5$ is the best value for both GPT-2 for Pythia.
Figure 5: The factual recall mechanism increases substantially across GPT-2 and Pythia after attention modification.
@@ Results.
We highlight the effect of our model editing method on the strength of the factual recall mechanism in Figure 5. Originally, GPT-2 has only 4% of the cases where the factual mechanism prevails the counterfactual one, and Pythia only 30%. However, after modifying the attention weights of the entries mentioned above, the strength of the factual mechanism increases drastically that it wins over the other mechanism in 50% of the cases for both models. This result is remarkable since we modify only two entries in the attention map out of the 33,264 attention values of GPT-2 (117M parameters) and three entries out of the 270,848 attention values of Pythia (6.9B parameters). This highlights the importance of the interpretability analysis in Sections 6.3 and 6.2, which enables us to find the detailed role played by the individual units of each transformer layer.
@@ 6.5 What Word Choices Intensify the Competition?
After the intrinsic intervention to edit the internal states of the model, we explore how the similarity between $t_{\mathrm{fact}}$ and $t_{\mathrm{cofa}}$ in our dataset affects the mechanism described in the previous sections.
Method.
We divided the dataset into 10 equal bins based on the similarity between the vectors for $t_{\mathrm{fact}}$ and $t_{\mathrm{cofa}}$, with each bin containing $1000$ items. Starting from the lowest, each group represents a $10\%$ segment of the dataset, arranged by increasing similarity scores. For our word similarity metric, we calculate the cosine similarity of the 300-dimensional word embeddings from the pre-trained Word2Vec model (Mikolov et al., 2013) implemented in the Gensim Python package (Řehůřek and Sojka, 2010).
Figure 6: Prediction frequency of factual token by similarity level. We show the percentage of $t_{\mathrm{fact}}$ predictions within each bin compared to the entire dataset (represented by a dotted line) across various model sizes. We can notice that more similar $t_{\mathrm{fact}}$ and $t_{\mathrm{cofa}}$ are, and the factual mechanism is stronger.
@@ Results.
As a result of the varying similarity of the two tokens, we see a drastic change in the dominance of the factual mechanism in Figure 6.
Finding 1: Similar tokens confuse the model more easily. Consistently across all the models, the more similar the two tokens are, the more likely the model is to be confused and mistakenly let the factual mechanism dominate, to predict $t_{\mathrm{fact}}$ as its output.
Finding 2: Larger models suffer more from such confusion. For example, the largest one, Pythia-6.9B, demonstrates a very strong attachment to the factual information, letting the factual mechanism win almost 45% of the cases when the token similarity reaches 90%. Even when the similarity is low, larger models are still more likely to confuse and lean towards the factual mechanism. This finding resonates with the observations from the inverse scaling prize McKenzie et al. (2023) that larger models have a greater capacity to store and retrieve factual information, thus more influenced by the factual mechanism.
@@ 7 Discussion and Future Work
Situating Our Findings in Related Work.
Our findings about the late attention blocks are consistent with Geva et al. (2023), showing that late attention blocks write most of the information to the last layer when adding a counterfactual premise. Surprisingly, however, we find that the largest contribution to the factual prediction of the network mostly comes from the suppression of the counterfactual token read from the attribute position rather than the promotion of the factual token from the subject position.
Consistently with McDougall et al. (2023), we find that few highly specialized heads suppress the counterfactual information. Moreover, we make a unique contribution up-weighting only two or three attention entries of these heads to increase substantially the number of factual responses of the model.
With an approach similar to ours, Yu et al. (2023) find that more heads can promote the factual mechanism, also in early layers, but found it challenging to improve the factual responses by scaling up the weights of the attention maps. This discrepancy can be due to the broader set of topics we include in our prompts, which allowed us to select fewer, more specialized heads, to the different ways the prompts are framed, or also to our more focused modification of the attention maps.
Future Work
For future research directions, we aim to analyze more in depth how our findings depend on the prompt structure and whether the promotion of factual responses by suppressing the counterfactuals generalizes to larger models and a more comprehensive variety of datasets.
@@ 8 Conclusion
In this work, we have proposed the formulation of the competition of mechanisms as a powerful interpretation when LLMs need to handle multiple mechanisms, only one of which leads to the correct answer. We deployed two mechanistic interpretability tools, logit inspection and attention modification, and identified critical positions and model components involved in competing for the mechanisms. Finally, we discovered a few localized positions in the attention map, which largely control the strength of the factual mechanism. Our study sheds light on future work on interpretability research for LLMs.
@@
Limited models: Our study aligns with most existing work in mechanistic interpretability to use GPT-2 small. However, we understand that this is a small model with far fewer parameters than current state-of-the-art LLMs. Future work is welcome to extend to larger-sized models, which might generalize our conclusion to a certain extent, and also reveal interesting behavior once the models get beyond a specific size, maybe also seeing a U-shaped curve Wei et al. (2023) for the dominance of the counterfactual mechanism.
Interpretability method: Furthermore, our experiments and insights are heavily grounded in the interpretability within the embedding space of the model’s inner components. This approach is reliable and extensively employed in mechanistic interpretability research (Dar et al., 2023; Geva et al., 2022; Halawi et al., 2023). The logit inspection method, although commonly employed in previous work, can occasionally fail to reflect the actual importance of some vocabulary items, especially in the early layers of the network (Belrose et al., 2023).
Simplicity of the prompts: Our prompts have a relatively simple structure for the controllability of the counterfactual information tracing, as it is very challenging to follow the information flow in a more diversified set of prompts. We welcome future work to explore methodological advances to enable analyses over more diverse prompts.
@@ Acknowledgment
This material is based in part upon works supported by the German Federal Ministry of Education and Research (BMBF): Tübingen AI Center, FKZ: 01IS18039B; by the Machine Learning Cluster of Excellence, EXC number 2064/1 – Project number 390727645; by the John Templeton Foundation (grant #61156); by a Responsible AI grant by the Haslerstiftung; and an ETH Grant (ETH-19 21-1). Alberto Cazzaniga and Diego Doimo are supported by the project “Supporto alla diagnosi di malattie rare tramite l’intelligenza artificiale” CUP: F53C22001770002. Alberto Cazzaniga received funding by the European Union – NextGenerationEU within the project PNRR ”PRP@CERIC” IR0000028 - Mission 4 Component 2 Investment 3.1 Action 3.1.1. Zhijing Jin is supported by PhD fellowships from the Future of Life Institute and Open Philanthropy.
@@
Appendix A Experiments for Pythia-6.9b
This section extends the experimental analysis conducted on GPT-2 to Pythia-6.9b. The goal is to replicate the prior methodology and compare the outcomes across the two different models, thus contributing to a broader understanding of model behaviors under similar conditions.
A.1 Macroscopic Inspection across Layers and Token Positions
Figure 7 provides a comparative analysis of the logit values for two specific tokens, labeled as factual and counterfactual, across various positions and layers in Pythia-6.9b.